# Planar graphs embedded in generic ways and realizing them as Reeb graphs of real algebraic functions

### v1

@@ Abstract.
Let $G$ be a graph admitting an M-generic embedding into ${\mathbb{R}}^{2}$. We have a real algebraic function on some compact and connected real algebraic manifold which is Morse-Bott and represented as the composition of some real algebraic map as in Theorem 2, presented later, with the canonical projection ${\pi}_{2,1}$, and whose Reeb graph is homeomorphic to $G$.
@@
We omit precise exposition on CW-complexes. A topological manifold has the structure of a CW complex. A smooth manifold has the structure of a CW complex canonically, which is a so-called PL manifold. Hereafter, unless otherwise stated, a smooth manifold is seen as a PL manifold when it is seen as a CW complex.
A piecewise smooth map $c:X\rightarrow Y$ on a CW complex $X$ into another CW complex $Y$ means a continuous map which is, for each cell $e_{X}$ of the CW complex and around each point $p_{e_{X}}\in e_{X}-F_{e_{X}}$, seen as a smooth map into some cell $e_{p,Y}$, on the complementary set of some set $F_{e_{X}}$ whose Lebesgue measure can be calculated as 0 of the cell $e_{X}$ of the CW complex of the domain. A piecewise smooth embedding means a piecewise smooth map which is also an embedding. A piecewise smooth homeomorphism is defined as a piecewise smooth map being also a homeomorphism and a piecewise smooth isotopy means an isotopy considered in the topology category which is also a piecewise smooth map.
An isomorphism between CW complexes means a piecewise smooth homeomorphism mapping each cell of a CW complex onto some cell in another CW complex.
@@
An isomorphism between two graphs is defined as a piecewise smooth homeomorphism mapping the vertex set of a graph onto that of the other. Two graphs are isomorphic if there exists an isomorphism between them.
@@
An embedding $e_{G}:G\rightarrow{\mathbb{R}}^{2}$ of a graph is generic with respect to ${\mathbb{R}}^{2}$ if the composition of this with the projection ${\pi}_{2,1}$ is a piecewise smooth embedding on each edge of the graph. The embedding is M-generic with respect to the projection if for every vertex $v$ of degree at least $2$ the image of a small open neighborhood $N(v)$ of $v$ by the composition of the embedding $e_{G}$ with the projection contains the value at $v$ in the interior of the image considered in $\mathbb{R}$.
@@
For an embedding $e_{G}:G\rightarrow{\mathbb{R}}^{2}$ of a graph $G$ being generic with respect to ${\mathbb{R}}^{2}$, there exists piecewise smooth isotopy $E_{G}:G\times\{t\mid 0\leq t\leq 1\}\rightarrow{\mathbb{R}}^{2}$ satisfying the following.
@@
    At each vertex $v$ of $G$, we have the values $E_{G}(v,t)$ ($0\leq t\leq 1$) as a constant value.
@@
    The image $E_{G}(G\times\{1\})$ is a union of straight segments in ${\mathbb{R}}^{2}$ where for any distinct pair of these sraight segments, they have exactly one point in common or do not intersect.
If we consider embeddings of graphs being M-generic with respect to the projection, then the restrictions of $E_{G}$ to the sets $e\times\{t\}$ ($0\leq t\leq 1$) are always M-generic with respect to the projection.
In short, a piecewise smooth isotopy of Proposition 1 preserves the values at vertices of the graphs of the functions represented as the compositions of the embeddings of the graphs with the projection and the genericity and M-genericity of the embeddings.
@@
Note also that the quotient spaces here are defined for (continuous) maps between general topological spaces and that we can enjoy similar arguments.
@@
The following is a kind of fundamental theorems in reconstructing real algebraic maps onto the closure of a given non-empty open set in a real affine space, first presented in [8]. This also extends an essential part of [7], in which the case of manifolds $S_{j}$ being mutually disjoint has been investigated. For [7], see also [9].
@@
      For each integer $j_{0}$$^{\prime}$ with $1\leq{j_{0}}^{\prime}\leq l^{\prime}$, there exists at least one integer $j_{0}$ satisfying $m_{l_{1}}(j_{0})={j_{0}}^{\prime}$. A non-negative integer $I_{j^{\prime}}$ is corresponded to each integer $1\leq j^{\prime}\leq l^{\prime}$.
@@
Note that originally, Theorem 2 is presented shown in a certain generalized situation. For example, in [15], regions surrounded by such circles and straight lines appear explicitly.
@@
The preimage of a sufficiently small open neighborhood diffeomorphic to $D^{2}$ is regarded as a copy of $D^{2}\times{\prod}_{j^{\prime}\in{\mathbb{N}}_{l^{\prime}}}S^{j^{\prime}}$. The map gives a smooth map with no singular point onto the $2$-dimensional disk, locally, for suitable local coordinates. Furthremore, the preimages of each point for the map are diffeomorphic to ${\prod}_{j^{\prime}\in{\mathbb{N}}_{l^{\prime}}}S^{j^{\prime}}$.
@@
We present two consecutive important propositions. We omit a precise presentation on the proof. For the proof, see also the method in the paper [2], which obtains a region in ${\mathbb{R}}^{2}$ from a given graph in ${\mathbb{R}}^{2}$ surrounded by mutually disjoint $1$-dimensional real algebraic manifolds and canonically collapsing to a given graph.
@@
We have a small regular neighborhood $N(E_{G}(G\times\{1\}))\subset{\mathbb{R}}^{2}$ of the graph $E_{G}(G\times\{1\})$ in the PL category. We have this in such a way that its boundary is a union of straight segments and homeomorphic to a disjoint union of copies of the circle $S^{1}$. We can also choose this in such a way that the restriction of ${\pi}_{2,1}$ to each connected component of the boundary of $N(E_{G}(G\times\{1\}))$ is generic with respect to ${\pi}_{2,1}$, where the circles are seen as graphs, canonically. This is not M-generic with respect to the projection, in general.
@@
    For each connected component of the boundary of $N(E_{G}(G\times\{1\}))$, seen as a graph, two adjacent edges $e_{v,1}$ and $e_{v,2}$ of it having exactly one vertex $v$ of it in common, the closure of one of these two edges of it is a straight segment in the straight line represented by the form $\{(t,y_{v})\mid t\in\mathbb{R}\}$ for a suitable number $y_{v}$.
@@
      The quotient space $Q_{{\pi}_{2,1}}(N(E_{G}(G\times\{1\})))$ is defined in the same way as that in defining the Reeb graph of a smooth function on a manifold with no boundary. Let $q_{{\pi}_{2,1},N(E_{G}(G\times\{1\}))}$ denote the quotient map.
@@
    The Poincaré-Reeb graph of $N(E_{G}(G\times\{1\}))$ collapses to a graph homeomorphic to $E_{G}(G\times\{1\})$. More precisely, for each vertex corresponding to the graph $G$ of degree greater than $3$, exactly one edge of the graph $Q_{{\pi}_{2,1}}(N(E_{G}(G\times\{1\})))$ (resp. $Q_{{\pi}_{S^{1}}}(N(E_{G}(G\times\{1\})))$) is corresponded to and by collapsing them, we have a graph homeomorphic to $G$ and $E_{G}(G\times\{1\})$.
@@
For each vertex $v_{E_{G}}$ of the graph $E_{G}(G\times\{1\}))$ which is not of degree $2$, we have a neighborhood $N(v_{E_{G}})\subset{\mathbb{R}}^{2}$ of $v_{E_{G}}$ which is diffeomorphic to $D^{2}$ and whose boundary is a circle centered at $v_{E_{G}}$ and of a fixed sufficiently small radius. We can choose the piecewise smooth isotopy $E_{G}$ and the set $N(E_{G}(G\times\{1\}))\subset{\mathbb{R}}^{2}$ in Propositions 2 and 3 in such a way that for each connected component $E_{v_{E_{G}}}$ of $N(v_{E_{G}})-(N(v_{E_{G}})\bigcap E_{G}(G\times\{1\}))$, exactly one vertex $v_{E_{v_{E_{G}}}}$ of degree $2$ of exactly one connected component $C_{E_{v_{E_{G}}}}$ of the boundary of $N(E_{G}(G\times\{1\}))$ where the value of the restriction of ${\pi}_{2,1}$ to the unique connected component $C_{E_{v_{E_{G}}}}$ of the boundary of $N(E_{G}(G\times\{1\}))$ is not in the image of a small open neighborhood $v_{E_{v_{E_{G}}}}\in U_{E_{v_{E_{G}}}}\subset C_{E_{v_{E_{G}}}}$. In addition, every vertex $v_{E_{v_{E_{G}}}}$ of degree $2$ of every connected component $C_{E_{v_{E_{G}}}}$ of the boundary of $N(E_{G}(G\times\{1\}))$ appears in this way. Thus we have a natural correspondence.
For this, important arguments of [13] are respected, where we do not assume related knowledge. Our related arguments here are regarded as a kind of variants of the original arguments.
@@
We can choose a circle $S_{j}:=S_{v_{E_{G}}}$ centered at a suitable point and of a suitable sufficiently small radius for each vertex $v_{E_{G}}$ of the graph $E_{G}(G\times\{1\}))$ in STEP 1 satisfying the following.
  • The circle $S_{v_{E_{G}}}$ bounds the disk $D_{v_{E_{G}}}$ containing the vertex $v_{E_{v_{E_{G}}}}$ of degree $2$ of the connected component $C_{E_{v_{E_{G}}}}$ in its interior.
  • The circle $S_{v_{E_{G}}}$ passes through a point of the straight line of the form $\{({\pi}_{2,1}(v_{E_{G}}),t)\mid t\in\mathbb{R}\}$ and at the point the tangent vector spaces of $S_{v_{E_{G}}}$ and the straight line, which are of $1$-dimensional, agree.
  • The set $S_{v_{E_{G}}}\bigcap N(E_{G}(G\times\{1\}))$ is diffeomorphic to $D^{1}$.
  • The circles $S_{v_{E_{G}}}$ are mutually disjoint.
  • The restriction of ${\pi}_{2,1}$ to the 1-dimensional manifold $S_{v_{E_{G}}}\bigcap N(E_{G}(G\times\{1\}))$ has exactly one singular point and the unique point is in its interior and of the form $({\pi}_{2,1}(v_{E_{G}}),y_{v_{E_{G}}})$.
@@
We can choose a circle $S_{j}$ centered at a suitable point and of a suitable sufficiently small radius as presented in FIGURE 2 for each vertex $v_{N(E_{G}(G\times\{1\}))}$ of degree $2$ of each connected component $C_{N(E_{G}(G\times\{1\}))}$ of the boundary of $N(E_{G}(G\times\{1\}))$ where the value of the restriction of ${\pi}_{2,1}$ to the connected component of the boundary of $N(E_{G}(G\times\{1\}))$ is in the image of a small open neighborhood $v_{N(E_{G}(G\times\{1\}))}\in U_{v_{N(E_{G}(G\times\{1\}))}}\subset C_{N(E_{G}(G\times\{1\}))}$. We can also choose this as a circle bounding the disk $D_{j}$ containing $v_{N(E_{G}(G\times\{1\}))}$ in the interior. We can also choose remaining circles $S_{j}$ centered at suitable points and of suitable sufficiently small radii as presented in FIGURE 2 to have our desired region $D$. We can choose all chosen circles $S_{j}$ here satisfying the following conditions.
@@
  • The restriction of ${\pi}_{2,1}$ to the $1$-dimensional manifold $S_{j}\bigcap N(E_{G}(G\times\{1\}))$ has no singular point for circles being not chosen as $S_{v_{E_{G}}}$ in the first case in the present step.
@@
Figure 1. The black straight objects in the figure above shows the graph $E_{G}(G\times\{1\})$ around a vertex $v_{E_{G}}$ of the graph $E_{G}(G\times\{1\}))$ of degree $5$ and the blue object there shows the boundary of $N(E_{G}(G\times\{1\}))$. In the figure below, the green circles show circles and the green dotted lines show straight lines of the form $\{(t,y_{0})\mid t\in\mathbb{R}\}$ passing through the centers of the circles and the straight line of the form $\{({\pi}_{2,1}(v_{E_{G}}),t)\mid t\in\mathbb{R}\}$ is depicted by the black straight line.
Figure 2. The former two figures here show circles $S_{j}$ for vertices $v_{N(E_{G}(G\times\{1\}))}$ of degree $2$ of connected components $C_{N(E_{G}(G\times\{1\}))}$ of the boundary of $N(E_{G}(G\times\{1\}))$ where the values of the restrictions of ${\pi}_{2,1}$ to the connected components of the boundary of $N(E_{G}(G\times\{1\}))$ are in the images of small open neighborhoods $v_{N(E_{G}(G\times\{1\}))}\in U_{v_{N(E_{G}(G\times\{1\}))}}\subset C_{N(E_{G}(G\times\{1\}))}$. The last figure shows an explicit case for ”the remaining circles $S_{j}$”. In the present figures, objects which are colored in black or shaded are for $N(E_{G}(G\times\{1\}))$. The green circles show circles and the green dotted lines show straight lines passing through the centers of the circles and relative locations.
@@
We have first investigated functions represented as the compositions of maps as in Theorem 2 with ${\pi}_{2,1}$ in [8]. [15] also shows this where we need a little effort and nned to change regions there, surrounded by circles centered at several points and of fixed radii and straight lines, into regions as in Theorem 2. There, Reeb graphs are explicitly presented with the structures of the graphs. For example, [15] investigates cases of trees.

### v2

@@ Abstract.
Let $G$ be a graph admitting an embedding into ${\mathbb{R}}^{2}$ which is M-generic with respect to ${\pi}_{2,1}$. We have a real algebraic function on some compact and connected real algebraic manifold which is Morse-Bott and represented as the composition of some real algebraic map as in Theorem 2, presented later, with the canonical projection ${\pi}_{2,1}$, and whose Reeb graph is homeomorphic to $G$.
@@
We omit precise exposition on fundamental terminologies and propositions on CW complexes.
A topological manifold has the structure of a CW complex. A smooth manifold has the structure of a CW complex canonically, which is a so-called PL manifold. Hereafter, unless otherwise stated, a smooth manifold is seen as a PL manifold when it is seen as a CW complex.
A piecewise smooth map $c:X\rightarrow Y$ on a CW complex $X$ into another CW complex $Y$ means a continuous map which is, for each cell $e_{X}$ of the CW complex and some subset $F_{e_{X}}\subset e_{X}$ whose Lebesgue measure can be calculated as 0 in $e_{X}$, at each point $p_{e_{X}}\in e_{X}-F_{e_{X}}$, seen as a smooth map into some cell $e_{p,Y}$. A piecewise smooth embedding means a piecewise smooth map being also a topological embedding. A piecewise smooth homeomorphism is defined as a piecewise smooth map being also a homeomorphism and a piecewise smooth isotopy means an isotopy considered in the topology category which is also a piecewise smooth map.
An isomorphism between CW complexes means a piecewise smooth homeomorphism mapping each cell of a CW complex onto some cell in another CW complex and two CW complexes are isomorphic if an isomorphism exists between them.
@@
For a union of cells of a CW complex, a regular neighborhood of it is defined. We do not explain its rigorous definition and in short, it is a neighborhood which is a closed set in the CW complex and collapses to the union.
@@
As a specific case of CW complexes, an isomorphism between two graphs is a piecewise smooth homeomorphism mapping the vertex set of a graph onto that of the other and two graphs are isomorphic if there exists an isomorphism between them.
@@
An embedding $e_{G}:G\rightarrow{\mathbb{R}}^{2}$ of a graph is generic with respect to ${\mathbb{R}}^{2}$ if the composition of this with the projection ${\pi}_{2,1}$ is a piecewise smooth embedding on each edge of the graph. The embedding is M-generic with respect to the projection if for every vertex $v$ of degree at least $2$ the image of some small open neighborhood $N(v)$ of $v$ by the composition of the embedding $e_{G}$ with the projection contains the value at $v$ in the interior of the image considered in $\mathbb{R}$.
@@
For an embedding $e_{G}:G\rightarrow{\mathbb{R}}^{2}$ of a graph $G$ being generic with respect to the projection ${\pi}_{2,1}$, there exists piecewise smooth isotopy $E_{G}:G\times\{t\mid 0\leq t\leq 1\}\rightarrow{\mathbb{R}}^{2}$ satisfying the following.
@@
    The restrictions of $E_{G}$ to the sets $e\times\{t\}$ ($0\leq t\leq 1$) are always generic with respect to the projection.
  3. (3)
    At each vertex $v$ of $G$, we have the values ${\pi}_{2,1}\circ E_{G}(v,t)$ ($0\leq t\leq 1$) as a constant value.
@@
    The image $E_{G}(G\times\{1\})$ is a union of straight segments in ${\mathbb{R}}^{2}$ where for any distinct pair of these straight segments, they have exactly one point in common or do not intersect.
As a specific case, in the case where an embedding $e_{G}$ is M-generic with respect to the projection, the restrictions of $E_{G}$ to the sets $e\times\{t\}$ ($0\leq t\leq 1$) are always M-generic with respect to the projection ${\pi}_{2,1}$.
In short, a piecewise smooth isotopy of Proposition 1 preserves the values at vertices of the graphs of the functions represented as the compositions of the embeddings of the graphs with the projection ${\pi}_{2,1}$ and the (M-)genericity of the embeddings.
@@
Note also that the quotient spaces here are defined for (continuous) maps between general topological spaces and that we can enjoy similar arguments around fundamental definitions.
@@
The following is a kind of fundamental theorems in reconstructing real algebraic maps onto the closure of a given non-empty open set in a real affine space, first presented in [8]. This also extends an essential part of [7], in which the case of manifolds $S_{j}$ being mutually disjoint has been investigated. For, see also a related preprint [9], remarking [7].
@@
      For each integer $j_{0}$$^{\prime}$ with $1\leq{j_{0}}^{\prime}\leq l^{\prime}$, there exists at least one integer $j_{0}$ satisfying $m_{l_{1}}(j_{0})={j_{0}}^{\prime}$. A positive integer $I_{j^{\prime}}$ is corresponded to each integer $1\leq j^{\prime}\leq l^{\prime}$.
@@
Note that originally, Theorem 2 is shown in a certain generalized situation. For example, in [15], regions surrounded by such circles and straight lines appear explicitly.
@@
The preimage of a sufficiently small open neighborhood diffeomorphic to $D^{2}$ is regarded as a copy of $D^{2}\times{\prod}_{j^{\prime}\in{\mathbb{N}}_{l^{\prime}}}S^{j^{\prime}}$. The map gives a smooth map with no singular point onto the $2$-dimensional disk, locally, for suitable local coordinates. Furthermore, the preimages of each point for the map are diffeomorphic to ${\prod}_{j^{\prime}\in{\mathbb{N}}_{l^{\prime}}}S^{j^{\prime}}$.
@@
We present two consecutive important propositions. We omit a precise presentation on the proof. For the proof, the method in the paper [2] is also respected. [2] obtains a region in ${\mathbb{R}}^{2}$ from a given graph in ${\mathbb{R}}^{2}$ surrounded by mutually disjoint $1$-dimensional real algebraic manifolds and canonically collapsing to a given graph.
@@
In Proposition 1, we have a small regular neighborhood $N(E_{G}(G\times\{1\}))\subset{\mathbb{R}}^{2}$ of the graph $E_{G}(G\times\{1\})$ where we discuss in the PL category. We have this in such a way that its boundary is a union of straight segments and homeomorphic to a disjoint union of copies of the circle $S^{1}$. We can also choose this in such a way that the restriction of ${\pi}_{2,1}$ to each connected component of the boundary of $N(E_{G}(G\times\{1\}))$ is generic with respect to ${\pi}_{2,1}$, where the circle is seen as a graph, canonically. This is not M-generic with respect to the projection.
@@
    For each connected component of the boundary of $N(E_{G}(G\times\{1\}))$, seen as a graph, two adjacent edges $e_{v,1}$ and $e_{v,2}$ of it having exactly one vertex $v$ of it in common, the closure of exactly one of these two edges of it is a straight segment in the straight line represented by the form $\{(t,y_{v})\mid t\in\mathbb{R}\}$ for a suitable number $y_{v}$ and that of the other is not.
@@
      The quotient space $Q_{{\pi}_{2,1}}(N(E_{G}(G\times\{1\})))$of $N(E_{G}(G\times\{1\}))$ is defined in the same way as that in defining the Reeb graph of a smooth function on a manifold with no boundary by using he restriction of ${\pi}_{2,1}$. Let $q_{{\pi}_{2,1},N(E_{G}(G\times\{1\}))}$ denote the quotient map.
@@
    The Poincaré-Reeb graph of $N(E_{G}(G\times\{1\}))$ collapses to a graph homeomorphic to $E_{G}(G\times\{1\})$. More precisely, for each vertex of graph $G$ of degree greater than $3$, exactly one edge of the graph $Q_{{\pi}_{2,1}}(N(E_{G}(G\times\{1\})))$ is corresponded to, this yields a one-to-one correspondence and by collapsing these edges of the graph, we have a graph homeomorphic to $G$ and $E_{G}(G\times\{1\})$.
@@
We can choose the piecewise smooth isotopy $E_{G}$ and the set $N(E_{G}(G\times\{1\}))\subset{\mathbb{R}}^{2}$ in Propositions 2 and 3 in such a way that the following hold.
  • For each vertex $v_{E_{G}}$ of the graph $E_{G}(G\times\{1\})$ which is not of degree $2$, we have a neighborhood $N(v_{E_{G}})\subset{\mathbb{R}}^{2}$ of $v_{E_{G}}$ which is diffeomorphic to $D^{2}$ and whose boundary is a circle centered at $v_{E_{G}}$ and of a fixed sufficiently small radius.
  • For each connected component $E_{v_{E_{G}}}$ of $N(v_{E_{G}})-(N(v_{E_{G}})\bigcap E_{G}(G\times\{1\}))$, there exists exactly one vertex $v_{E_{v_{E_{G}}}}$ of degree $2$ of a connected component $C_{E_{v_{E_{G}}}}$ of the boundary of $N(E_{G}(G\times\{1\}))$ where the value of the restriction of ${\pi}_{2,1}$ to the connected component $C_{E_{v_{E_{G}}}}$ of the boundary of $N(E_{G}(G\times\{1\}))$ is not in the interior of the image of any small open neighborhood $v_{E_{v_{E_{G}}}}\in U_{E_{v_{E_{G}}}}\subset C_{E_{v_{E_{G}}}}$ of $v_{E_{v_{E_{G}}}}$ considered in $C_{E_{v_{E_{G}}}}$.
  • In addition, every vertex $v_{E_{v_{E_{G}}}}$ of degree $2$ of any connected component $C_{E_{v_{E_{G}}}}$ of the boundary of $N(E_{G}(G\times\{1\}))$ as above, or a vertex of $C_{E_{v_{E_{G}}}}$ where the value of the restriction of ${\pi}_{2,1}$ to a connected component $C_{E_{v_{E_{G}}}}$ of the boundary of $N(E_{G}(G\times\{1\}))$ is not in the interior of the image of any small open neighborhood $v_{E_{v_{E_{G}}}}\in U_{E_{v_{E_{G}}}}\subset C_{E_{v_{E_{G}}}}$ of $v_{E_{v_{E_{G}}}}$ considered in $C_{E_{v_{E_{G}}}}$, appears in this way. Thus we have a natural one-to-one correspondence. For this, important arguments of [13] are respected, where we do not assume related knowledge. Our related arguments here are regarded as a kind of variants of the original arguments.
@@
We can choose a circle $S_{j}:=S_{v_{E_{v_{E_{G}}}}}$ centered at a suitable point and of a suitable sufficiently small radius for each vertex $v_{E_{G}}$ of the graph $E_{G}(G\times\{1\})$ and each connected component $C_{E_{v_{E_{G}}}}$ of $N(v_{E_{G}})-(N(v_{E_{G}})\bigcap E_{G}(G\times\{1\}))$ with $v_{E_{v_{E_{G}}}}\in C_{E_{v_{E_{G}}}}$ in STEP 1 and satisfying the following.
  • The circle $S_{v_{E_{v_{E_{G}}}}}$ bounds the disk $D_{v_{E_{v_{E_{G}}}}}$ containing the vertex $v_{E_{v_{E_{G}}}}$ of degree $2$ of the connected component $C_{E_{v_{E_{G}}}}$ in its interior.
  • In the case the vertex $v_{E_{G}}$ of the graph $E_{G}(G\times\{1\})$ is of degree at least $3$, the circle $S_{v_{E_{v_{E_{G}}}}}$ passes through a point of the straight line of the form $\{({\pi}_{2,1}(v_{E_{G}}),t)\mid t\in\mathbb{R}\}$ and at the point the tangent vector spaces of $S_{v_{E_{v_{E_{G}}}}}$ and the straight line, which are of $1$-dimensional, agree.
  • The set $S_{v_{E_{v_{E_{G}}}}}\bigcap N(E_{G}(G\times\{1\}))$ is diffeomorphic to $D^{1}$.
  • The disks $D_{v_{E_{v_{E_{G}}}}}$ are mutually disjoint.
  • In the case the vertex $v_{E_{G}}$ of the graph $E_{G}(G\times\{1\})$ is of degree at least $3$, the restriction of ${\pi}_{2,1}$ to the 1-dimensional manifold $S_{v_{E_{v_{E_{G}}}}}\bigcap N(E_{G}(G\times\{1\}))$ has exactly one singular point and the unique point is in its interior and of the form $({\pi}_{2,1}(v_{E_{G}}),y_{v_{E_{v_{E_{G}}}}})$. In the case the vertex $v_{E_{G}}$ of the graph $E_{G}(G\times\{1\})$ is of degree $1$, the restriction of ${\pi}_{2,1}$ to the 1-dimensional manifold $S_{v_{E_{v_{E_{G}}}}}\bigcap N(E_{G}(G\times\{1\}))$ has no singular point.
@@
We can choose a circle $S_{j}$ centered at a suitable point and of a suitable sufficiently small radius as presented in FIGURE 2 for each vertex $v_{N(E_{G}(G\times\{1\}))}$ of $2$ of each connected component $C_{N(E_{G}(G\times\{1\}))}$ of the boundary of $N(E_{G}(G\times\{1\}))$ where the value of the restriction of ${\pi}_{2,1}$ to the connected component of the boundary of $N(E_{G}(G\times\{1\}))$ is in the interior of the image of a small open neighborhood $v_{N(E_{G}(G\times\{1\}))}\in U_{v_{N(E_{G}(G\times\{1\}))}}\subset C_{N(E_{G}(G\times\{1\}))}$ considered in $C_{N(E_{G}(G\times\{1\}))}$. We can also choose this as a circle bounding the disk $D_{j}$ containing $v_{N(E_{G}(G\times\{1\}))}$ in the interior.
We can also choose remaining circles $S_{j}$ centered at suitable points and of suitable sufficiently small radii as presented in FIGURE 2 to have our desired region $D$. For the last circle $S_{j}$, we use $D_{j}$ for the disk bounded by $S_{j}$. We can choose all chosen circles $S_{j}$ here satisfying the following conditions.
@@
  • The restriction of ${\pi}_{2,1}$ to the $1$-dimensional manifold $S_{j}\bigcap N(E_{G}(G\times\{1\}))$ has no singular point for circles being not chosen as $S_{v_{E_{v_{E_{G}}}}}$ in the present step.
  • The disks $D_{j}$ cover the boundary of of $N(E_{G}(G\times\{1\}))$.
@@
Figure 1. The black straight objects in the figure above show the graph $E_{G}(G\times\{1\})$ around a vertex $v_{E_{G}}$ of the graph $E_{G}(G\times\{1\})$ of degree $5$, the purple straight objects show the graph $Q_{{\pi}_{2,1}}(N(E_{G}(G\times\{1\})))$ locally, and the blue objects show the boundary of $N(E_{G}(G\times\{1\}))$, which is also a graph. In the figure below, the green circles show important circles $S_{j}:=S_{v_{E_{v_{E_{G}}}}}$ and the green dotted lines show straight lines of the form $\{(t,y_{0})\mid t\in\mathbb{R}\}$ passing through the centers of the circles, each green dot shows a point in a green circle and the dotted green straight line associated to the circle and outside $N(E_{G}(G\times\{1\}))$. The straight lines of the form $\{({\pi}_{2,1}(v_{E_{G}}),t)\mid t\in\mathbb{R}\}$ are depicted by the black straight lines. Such objects also show relative locations of the shapes.
Figure 2. The former two figures here show circles $S_{j}$ for vertices $v_{N(E_{G}(G\times\{1\}))}$ of degree $2$ of connected components $C_{N(E_{G}(G\times\{1\}))}$ of the boundary of $N(E_{G}(G\times\{1\}))$ where the values of the restrictions of ${\pi}_{2,1}$ to the connected components of the boundary of $N(E_{G}(G\times\{1\}))$ are in the interiors of the images of some small open neighborhoods $v_{N(E_{G}(G\times\{1\}))}\in U_{v_{N(E_{G}(G\times\{1\}))}}\subset C_{N(E_{G}(G\times\{1\}))}$. The last figure shows an explicit case for ”the remaining circles $S_{j}$”. In these present figures, objects which are colored in black or shaded are for $N(E_{G}(G\times\{1\}))$. The green circles show important circles $S_{j}$ and the green dotted lines show straight lines of the form $\{(t,y_{0})\mid t\in\mathbb{R}\}$ passing through the centers of the circles. Such objects show relative locations of the shapes.
@@
We have first investigated functions represented as the compositions of maps as in Theorem 2 with ${\pi}_{2,1}$ in [8]. [15] also shows this where we need a little effort and need to change regions there, surrounded by circles centered at several points and of fixed radii and straight lines, into regions as in Theorem 2. There, Reeb graphs are explicitly presented with the structures of the graphs. For example, [15] investigates cases of trees.